# KPC-cF: Aspect-Based Sentiment Analysis via Implicit-Feature Alignment with Corpus Filtering

### v5

@@ 4.1 Dataset for Fine-Tuning and Test
Table 1: KR3 test set results for Aspect Category Detection (middle) and Aspect Category Polarity (right). We reported the computational requirements as #Training. Specifically, for Baseline+PL, it refers to the overall capacity of structured NLI data and total number of valid tokens from KR3 Train. The performances where filtered models outperform non-filtered models under the same training conditions are highlighted in bold & underline. KPC-cF†, is jointly fine-tuned on the Kor-SemEval & Filtered KR3 Train. Task Type is represented by Section§3.4. For GPT-4o mini, to reflect diversity in few-shot, we report the average of two results measured with temperature 0 or 0.7 and top-p 1, extending Wang et al. (2023). Zero-Shot, Trans-Ta (Li et al., 2020; Zhang et al., 2021) were evaluated using our training settings in Appendix§C.
@@ 7 Limitations and Future Work
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Table 7, 7). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None. Furthermore, our threshold filtering assigns a high-confidence Target None label for connotations with high similarity but lower errors, which supports a better distinction between semantic similarity and causal relationships. This is particularly effective when excluding personalized terms and Korean proper nouns, such as ‘Fixture (i.e., local native)’ used by reviewers. Our current translation-based training does not sufficiently capture the tone of sentences composed of proper nouns and particles from various domains within Korea. However, we have confirmed through various evaluations (Table 1, 2, 3, 4, 5, 7 and Figure 5) that our filtering approach aids in learning actual level of entailment relationships, moving beyond merely identifying similar relationships between sentences. To mitigate spurious correlations (Haig, 2003; Udomcharoenchaikit et al., 2022), we propose upper filtering that enables Target None predictions to function as data facilitating the model’s learning of true entailment relationships between sentences, without relying solely on surface-level similarity in NLI pairs.
@@ Acknowledgements
We thank Chanseo Nam (Korea National University of Education) for manually correcting and annotating the Kor-SemEval and KR3 testset used to test our model, during his undergraduate research internship. We also thank Kyeongpil Kang (Kangwon National University) for his valuable discussions. This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all.
Appendix A Mathematical Formulation under Assumption: Enhanced NLI-Based Pseudo-Classifier
A.1 Similarity and Confidence
@@
A.2 Relationship with Sementic Label Accuracy
@@
A.3 Loss Function and Gradients
@@
A.4 Updates to Model Parameters
@@
A.5 Gradient Propagation to Tokens
@@
A.6 Alignment and Overall Dynamics
@@
A.7 Extension of Experimental Results
@@
Figure 5: Performance of ACD and ACP on KR3 test set during adaptive fine-tuning ($D_{T}$ and ${D_{S}}_{t}\rightarrow D_{T}$). Left: results with the addition of other fine-tuned BaselineXLM-R. th denotes the threshold for confidence of pseudo-labeling, and L denotes the threshold for LaBSE filtering; Right: BaselineXLM-R tuning compared in this paper. Blue line represents model, KPC-cF.
@@ D.2 Preliminary Results
Appendix E Computational Resources

### v6

@@ 4.1 Dataset for Fine-Tuning and Test
Table 1: KR3 test set results for Aspect Category Detection (middle) and Aspect Category Polarity (right). We reported the computational requirements as #Training. Specifically, for Baseline+PL, it refers to the overall capacity of structured NLI data and total number of valid tokens from KR3 Train. The performances where filtered models outperform non-filtered models under the same training conditions are highlighted in bold & underline. KPC-cF†, is jointly fine-tuned on the Kor-SemEval & Filtered KR3 Train. Task Type is represented by Section§3.4. For GPT-4o mini, Babel-9B-Chat, to reflect diversity in few-shot, we report the average of two results measured with temperature 0 or 0.7 and top-p 1, extending Wang et al. (2023). Zero-Shot, Trans-Ta (Li et al., 2020; Zhang et al., 2021) were evaluated using our settings in Appendix§C.
@@ 7 Limitations and Future Work
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Table 7, 7). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None. Furthermore, our threshold filtering assigns a high-confidence Target None label for connotations with high similarity but lower errors, which supports a better distinction between semantic similarity and causal relationships. This is particularly effective when excluding personalized terms and Korean proper nouns, such as ‘Fixture (i.e., local native)’ used by reviewers. Our current translation-based training does not sufficiently capture the tone of sentences composed of proper nouns and particles from various domains within Korea. However, we have confirmed through various evaluations (Table 1, 2, 3, 4, 5, 7, 11 and Figure 5, 6, 7, 8) that our filtering approach aids in learning actual level of entailment relationships, moving beyond merely identifying similar relationships between sentences. To mitigate spurious correlations (Haig, 2003; Udomcharoenchaikit et al., 2022), we propose upper filtering that enables Target predictions to function as data facilitating the model’s learning of true entailment relationships between sentences, without relying solely on surface-level similarity in NLI pairs.
@@ Acknowledgements
We thank Chanseo Nam for manually correcting and annotating the Kor-SemEval and KR3 testset used to test our model, during his undergraduate research. We also thank Kyeongpil Kang, Byungok Han, Woohan Yun, Soyoung Yang, Joosun Yum and Guijin Son for their valuable discussions. Thanks to DAVIAN, MODULABS and UST AI as well. This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all.
Appendix A Formulation under Assumption: Enhanced NLI-Based Pseudo-Classifier
Similarity and Confidence
@@ Appendix A Formulation under Assumption: Enhanced NLI-Based Pseudo-Classifier
Relationship with Sementic Label Accuracy
@@
Loss Function and Gradients
@@
Updates to Model Parameters
@@
Gradient Propagation to Tokens
@@
Alignment and Overall Dynamics
@@
Extension of Experimental Results
@@
Figure 5: Performance of ACD and ACP on KR3 test set during adaptive fine-tuning ($D_{T}$ and ${D_{S}}_{t}\rightarrow D_{T}$). Left: results with the addition of other fine-tuned BaselineXLM-R. th denotes the threshold for confidence of pseudo-labeling, and L denotes the threshold for LaBSE filtering; Right: BaselineXLM-R tuning compared in this paper. Blue line represents KPC-cF.
A.1 Gradient Stability and Directionality from Dual Scoring
Notation. Let $x_{s},x_{a}\in\mathbb{R}^{d}$ denote the sentence embeddings in a compact metric space $\mathcal{E}$. Let $\Phi(x_{s},x_{a})\in\mathcal{E}\subset\mathbb{R}^{d}$ be the fused input representation. Let $\theta=\{W_{H},W_{V},W_{KQ}\}$ be the trainable parameters. Define:
$\text{sim}(x_{s},x_{a})$ $=\frac{\langle x_{s},x_{a}\rangle}{\|x_{s}\|\cdot\|x_{a}\|}\,\in[-1,1],$
$C(x_{s},x_{a})$ $=\max_{k}\,p(k\mid x_{s},x_{a};\theta)\,\in[0,1],$
$\mathcal{D}_{f}$ =$=\left\{(x_{s},x_{a},\hat{y})\in\mathcal{D}\,\middle|\,\text{sim}(x_{s},x_{a})>\tau_{1},\;C(x_{s},x_{a})>0.5\right\}$
where $\tau_{1}=\mathbb{E}_{\mathcal{D}}[\text{sim}]$ is the empirical mean similarity. The fixed confidence threshold of $0.5$ reflects the theoretical minimum for non-random predictions and ensures robustness under calibration skewness.
Let the empirical loss be:
$\mathcal{L}(\theta)=\frac{1}{|\mathcal{D}_{f}|}\sum_{(x_{s},x_{a},\hat{y})\in\mathcal{D}_{f}}\ell(f_{\theta}(x_{s},x_{a}),\hat{y})$
where $\ell$ is the cross-entropy loss.
Assumptions. (1) $\mathcal{E}$ is compact. (2) sim and $C$ are Lipschitz-continuous. (3) $\ell$ is locally convex and differentiable.
Lemma 1. Gradient Stability via Score Concentration
Let $L$ be the Lipschitz constant of sim.
$|\text{sim}(x_{1},x_{2})-\text{sim}(x_{1},x_{3})|\leq L\cdot\|x_{2}-x_{3}\|_{2}$
After filtering, we observe:
$\mu_{\text{sim}}^{f}$ $:=\mathbb{E}_{\mathcal{D}_{f}}[\text{sim}]>\mathbb{E}_{\mathcal{D}}[\text{sim}]=:\mu_{\text{sim}},$
$\sigma_{\text{sim}}^{f}$ $:=\sqrt{\mathbb{E}_{\mathcal{D}_{f}}[(\text{sim}-\mu_{\text{sim}}^{f})^{2}]}<\sqrt{\mathbb{E}_{\mathcal{D}}[(\text{sim}-\mu_{\text{sim}})^{2}]}=:\sigma_{\text{sim}},$
$\mu_{C}^{f}$ $:=\mathbb{E}_{\mathcal{D}_{f}}[C]>\mathbb{E}_{\mathcal{D}}[C]=:\mu_{C},$
$\sigma_{C}^{f}$ $:=\sqrt{\mathbb{E}_{\mathcal{D}_{f}}[(C-\mu_{C}^{f})^{2}]}<\sqrt{\mathbb{E}_{\mathcal{D}}[(C-\mu_{C})^{2}]}=:\sigma_{C}$
Define filtered gradient:
$\nabla_{\theta}\mathcal{L}_{f}=\frac{1}{|\mathcal{D}_{f}|}\sum_{(x_{s},x_{a},\hat{y})\in\mathcal{D}_{f}}\nabla_{\theta}\ell(f_{\theta}(x_{s},x_{a}),\hat{y})$
Score concentration implies bounded gradient deviation:
$\forall(x_{s},x_{a})\in\mathcal{D}_{f},\quad\mathbb{E}\left[\|\nabla_{\theta}\ell(f_{\theta}(x_{s},x_{a}),\hat{y})-\nabla_{\theta}\mathcal{L}_{f}\|_{2}^{2}\right]\leq\varepsilon_{1}(\sigma_{\text{sim}}^{f},\sigma_{C}^{f})$
Lemma 2. Confidence-Guided Gradient Direction under Reinitialization
Let the classifier weights be initialized as $W_{H}^{(t_{0})}\sim\mathcal{N}(0,\sigma^{2})$, and let $(x_{s},x_{a},y)\in\mathcal{D}_{f}$ be a filtered sample with high confidence $C(x_{s},x_{a})>\tau$. Then, at the training step $t=t_{0}$, where the classification head is reinitialized, the gradient satisfies:
$\left\langle\nabla_{W_{H}}\ell(f_{t_{0}}(x_{s},x_{a}),y),\Phi(x_{s},x_{a})\right\rangle>0$
indicating that gradient updates are initially aligned with the semantic direction $\Phi(x_{s},x_{a})$. This alignment facilitates faster convergence during early fine-tuning.
Lemma 3. Compactness and Gradient Directionality
Let $\{\Phi(x_{s}^{t},x_{a}^{t})\}_{t=1}^{\infty}\subset\mathcal{E}$ be a sequence of fused representations, where $\mathcal{E}$ is the compact representation space defined in Assumption (1).
By the topological definition of compactness, for every open cover $\{U_{\alpha}\}_{\alpha\in A}$ of $\mathcal{E}$, there exists a finite subcover. Hence, the sequence $\{\Phi(x_{s}^{t},x_{a}^{t})\}$ cannot diverge or oscillate unboundedly, and must lie within a finite union of open sets. This guarantees the existence of a convergent subsequence:
$\Phi(x_{s}^{t_{k}},x_{a}^{t_{k}})\rightarrow\Phi^{*}\in\mathcal{E}\quad\text{as}\quad k\rightarrow\infty.$
Assuming $\ell$ is continuous in $\Phi$ and $f_{\theta}$ is differentiable, we obtain:
$\nabla_{\theta}\ell(f_{\theta}(\Phi(x_{s}^{t_{k}},x_{a}^{t_{k}})),y_{t_{k}})\rightarrow\nabla_{\theta}\ell(f_{\theta}(\Phi^{*}),y^{*}).$
Furthermore, let $q^{(t)}:=W_{Q}\Phi(x_{s}^{t},x_{a}^{t})$ and $k^{(t)}:=W_{K}\Phi(x_{s}^{t},x_{a}^{t})$ denote the query and key projections. By linearity and the convergence $\Phi(x_{s}^{t},x_{a}^{t})\to\Phi^{*}$, we have:
$\langle q^{(t)},k^{(t)}\rangle=\langle W_{Q}\Phi(x_{s}^{t},x_{a}^{t}),W_{K}\Phi(x_{s}^{t},x_{a}^{t})\rangle\rightarrow\langle W_{Q}\Phi^{*},W_{K}\Phi^{*}\rangle.$
Therefore, the attention score matrix defined by
$A^{(t)}:=\text{softmax}\left(\frac{q^{(t)}(k^{(t)})^{\top}}{\sqrt{d_{k}}}\right)$
converges to
$A^{*}:=\text{softmax}\left(\frac{W_{Q}\Phi^{*}(W_{K}\Phi^{*})^{\top}}{\sqrt{d_{k}}}\right),$
and the attention structure becomes stable across samples.
Finally, the gradient w.r.t. $W_{KQ}$ also converges:
$\nabla_{W_{KQ}}\ell(f_{\theta}(\Phi(x_{s}^{t},x_{a}^{t})),y_{t})\rightarrow\nabla_{W_{KQ}}\ell(f_{\theta}(\Phi^{*}),y^{*}),$
with directional alignment preserved:
$\left\langle\nabla_{W_{KQ}}\mathcal{L}_{f},\;W_{KQ}\Phi^{*}\right\rangle>0.$
Lemma 4. Logit Monotonicity Post-Filtering
Let $\sigma$ denote softmax and define logit scores:
$P_{\text{pre}}(y|x_{s},x_{a})$ $=\sigma(W_{H}^{(t_{0})\top}W_{V}^{(t_{0})}\Phi(x_{s},x_{a}))_{y},$
$P_{\text{post}}(y|x_{s},x_{a})$ $=\sigma(W_{H}^{(t)\top}W_{V}^{(t)}\Phi(x_{s},x_{a}))_{y}$
Then:
$P_{\text{post}}(y|x_{s},x_{a})-P_{\text{pre}}(y|x_{s},x_{a})>\delta>0$
indicating alignment with Assumption 3.2.
Theorem. Convergence under Low Gradient Dispersion
Assume:
$\forall(x_{s},x_{a})\in\mathcal{D}_{f},\quad\mathbb{E}\left[\|\nabla_{\theta}\mathcal{L}-\mathbb{E}[\nabla_{\theta}\mathcal{L}]\|^{2}\right]\leq\sigma^{2}$
Then for $L$-smooth $\mathcal{L}$ and Adam optimizer:
$\lim_{t\to\infty}\left\|\frac{1}{|\mathcal{D}_{f}|}\sum_{(x_{s},x_{a},\hat{y})}\nabla_{\theta}\ell(f_{\theta}(x_{s},x_{a}),\hat{y})\right\|_{2}\to 0$
Summary of Theoretical Guarantees
The dual-filtered set $\mathcal{D}_{f}$ satisfies:
[…]
Under the NTK regime (Jacot et al., 2018), these enhancements extend to shallow-layer generalization (Table 4), improving decision boundary sharpness.
Figure 6: Multi-Head Attention Visualization. In intermediate layers associated with domain and linguistic understanding (Clark et al., 2019), the proposed KPC-cF model exhibits stronger self-attention activations across heads for tokens relevant to target language relations. In the final encoder layer responsible for classification, KPC-cF shows greater focus on label-relevant information in the [CLS] embedding, with reduced influence from irrelevant noise. This visualization is consistent with Lemma Lemma 3. Compactness and Gradient Directionality.
@@ D.2 Preliminary Results
Appendix E Attention Head Visualization
For token visualization (Figure 6, 7, 8), we used input sentences from the KR3 dataset, structured as [CLS] review [SEP] aspect [SEP]. An example snippet in Figure 6 is as follows: review = “맛 집이라고 찾아서 갔는데 손님이 오든 말든 불친절에 물컵에 기름기가 잘잘 흐르고 위생도 꽝이고 면에서는 밀가루 냄새만 나고 먹다가 다 남기고 실망하고 나왔어요. 왜 이런 집이 맛 집으로 평가받는지 모르겠다.”, aspect = “음식”.
Appendix F Computational Resources
@@
Model GPU.t Thruput PL Filtering PL Alignment Aspect Category Polarity
LaBSEavg MSP0.5 METEOR SemScore Precision Recall Micro-F1 4-way acc 3-way acc Binary
BaselineXLM-R+TR+PL 41 30.45K 48.93 65.85 91.5 84.23 87.69 84.21 86.57 90.35
BaselineXLM-R+TR+PL 31 22.00K 48.84 66.62 91.63 85.51 88.44 83.31 85.66 89.78
BaselineXLM-R+TR+PL 38 29.53K 48.99 65.75 90.39 84.18 87.14 84.50 86.88 90.86
KPC-cF 29 21.30K 49.03 66.58 92.46 84.53 88.29 84.34 86.72 90.89
Table 11: Ablation Study LaBSE-based similarity thresholding improves aspect detection, while confidence-based filtering via MSP enhances polarity classification. The proposed KPC-cF achieves complementary effects by combining both filters, yielding robust performance with reduced data. Throughput denotes the total number of training samples processed across both pre- and second-stage. The joint-trained KPC-cF † required 47 minutes of GPU time. Evaluation of PL sample quality using METEOR and SemScore (Aynetdinov & Akbik, 2024). Dual-filtered samples in KPC-cF show better input-label alignment than unfiltered PL data.
Figure 7: Multi-Head Attention Visualization. In intermediate layers associated with domain and linguistic understanding (Clark et al., 2019), the proposed KPC-cF model exhibits stronger self-attention activations across heads for tokens relevant to target language relations. In the final encoder layer responsible for classification, KPC-cF shows greater focus on label-relevant information in the [CLS] embedding, with reduced influence from irrelevant noise. This visualization is consistent with Lemma Lemma 3. Compactness and Gradient Directionality. Snippet is as follows: review = “맛 집이라고 찾아서 갔는데 손님이 오든 말든 불친절에 물컵에 기름기가 잘잘 흐르고 위생도 꽝이고 면에서는 밀가루 냄새만 나고 먹다가 다 남기고 실망하고 나왔어요. 왜 이런 집이 맛 집으로 평가받는지 모르겠다.”, aspect = “서비스”.
Figure 8: Multi-Head Attention Visualization. In intermediate layers associated with domain and linguistic understanding (Clark et al., 2019), the proposed KPC-cF model exhibits stronger self-attention activations across heads for tokens relevant to target language relations. In the final encoder layer responsible for classification, KPC-cF shows greater focus on label-relevant information in the [CLS] embedding, with reduced influence from irrelevant noise. This visualization is consistent with Lemma Lemma 3. Compactness and Gradient Directionality. Snippet is as follows: review = “편안한 느낌의 토스트 킴. 식사 한 끼로 좋은 토스트를 팝니다. 고기를 좋아해서 갈 때마다 시키는 건 불고기, 삼겹살, 돈가스, 소고기 치즈 토스트 등… 각자의 매력이 있고 한국식 토스트인데 정말 두둑합니다… 단품 세트로 시키면 감자튀김에 음료도 나오니 웬만한 패스트푸드 보다 낫다 보시면 되겠네요… 오천 원 하나로 두둑이 먹을 수 있어요.”, aspect = “가격”.